OpenStudy (anonymous):

Solve the inequality: (x + 3)(x - 5) < 0. A. x < -3 or x > 5 B. x < -5 or x > 3 C. -3 < x < 5 D. -5 < x < 3

OpenStudy (anonymous):

c

OpenStudy (anonymous):

what about this? Solve the inequality: (x + 3)(x - 5) > 0. A. x < -3 or x > 5 B. x < -5 or x > 3 C. -3 < x < 5 D. -5 < x < 3

OpenStudy (anonymous):

a

OpenStudy (anonymous):

the sign is flipped?

OpenStudy (anonymous):

you have to solve for x first

OpenStudy (anonymous):

-3,5

OpenStudy (anonymous):

\(\Large (x + 3)(x - 5) < 0\) \(\Large \color{red}{Case \ 1:} \begin{cases} x+3<0 \\ x-5>0 \end{cases}\Leftrightarrow \begin{cases} x<-3 \\ x>5 \end{cases}\) \(\Large \Rightarrow \)No solution \(\Large \color{red}{Case \ 2:}\begin{cases} x+3>0 \\ x-5<0 \end{cases}\Leftrightarrow \begin{cases} x>-3 \\ x<5 \end{cases} \Rightarrow -3<x<5\) Thus, the corect answer is \(\huge \color {blue}{C}\)

OpenStudy (anonymous):

look how @JungHyunRan broke it down for you

OpenStudy (anonymous):

ah got it, thanks